OpenStudy (anonymous):

Solve the Equation 4d+2d-4d+4=3d A.) 4/7 B.) 4 C.) -4/5 D.) -4

OpenStudy (anonymous):

it is B...

OpenStudy (anonymous):

4d+2d-4d+4=3d Simple and best practice solution for 4d+2d-4d+4=3d equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 4d+2d-4d+4=3d equation: Simplifying 4d + 2d + -4d + 4 = 3d Reorder the terms: 4 + 4d + 2d + -4d = 3d Combine like terms: 4d + 2d = 6d 4 + 6d + -4d = 3d Combine like terms: 6d + -4d = 2d 4 + 2d = 3d Solving 4 + 2d = 3d Solving for variable 'd'. Move all terms containing d to the left, all other terms to the right. Add '-3d' to each side of the equation. 4 + 2d + -3d = 3d + -3d Combine like terms: 2d + -3d = -1d 4 + -1d = 3d + -3d Combine like terms: 3d + -3d = 0 4 + -1d = 0 Add '-4' to each side of the equation. 4 + -4 + -1d = 0 + -4 Combine like terms: 4 + -4 = 0 0 + -1d = 0 + -4 -1d = 0 + -4 Combine like terms: 0 + -4 = -4 -1d = -4 Divide each side by '-1'. d = 4 Simplifying d = 4